supply price = 50+q/2 demand price = 150-q/5 find producer and consumer surplus


1
Expert's answer
2021-03-30T07:52:51-0400
"q_s=2p-100"

"q_d=750-5p"

"2p-100=750-5p"

"p=121"

"q=145"

"CS=\frac{145\times (150-121)}{2}=2,102.5"

"PS=\frac {120-50}{2}\times145=5,057"